# Review of Generative AI Methods in Cybersecurity

### 13 Mar 2024 v1

@@ Abstract
Large language models (LLMs) and generative artificial intelligence (GenAI) constitute paradigm shifts in cybersecurity that present hitherto unseen challenges as well as opportunities. In examining the state-of-the-art application of GenAI in cybersecurity, this work highlights how models like Google’s Gemini and ChatGPT-4 potentially enhance security protocols, vulnerability assessment, and threat identification. Our research highlights the significance of a novel approach that employs LLMs to identify and eliminate sophisticated cyber threats. This paper presents a thorough assessment of LLMs’ ability to produce important security insights, hence broadening the potential applications of AI-driven cybersecurity solutions. Our findings demonstrate the significance of GenAI in improving digital security. It offers recommendations for further investigations into the intricate relationship between cybersecurity requirements and artificial intelligence’s potential.
@@ 1 Introduction
The emergence of Generative Artificial Intelligence (GenAI) is heralding a revolutionary era in digital technology, particularly in cybersecurity and multi-modal interactions. Large Language Models (LLMs) and Natural Language Processing (NLP) advancements are leading the way, redefining AI’s capabilities as evidenced by the emergence of models like Google’s Gemini and ChatGPT. Improvements in security vulnerability scanning apps and deeper vulnerability analysis made possible by these developments surpass the capabilities of traditional Static Application Security Testing (SAST) techniques [1].
@@
Language models are essential in many sectors, including commerce, healthcare, and cybersecurity. Their progress shows a definite path from basic statistical methods to sophisticated neural networks [2], [3]. NLP skill development has benefited immensely from the use of LLMs. However, despite these advancements, a number of issues remain, including moral quandaries, the requirement to reduce error rates, and making sure that these models are consistent with our moral values. To solve these issues, moral monitoring and ongoing development are required.
2 The Challenges of GenAI
@@
Mohammed et al [4] define key challenges of the use of ChatGPT in cybersecurity, including: Exploring the Impact of ChatGPT on Cybersecurity; Constructing Honeypots; Code Security Enhancement; Misuse in Malware Development; Vulnerability Exploration; Propagating Disinformation; Attacks on Industrial Systems; Changing the Cyber Threat Landscape; Adapting Cybersecurity Strategies; and Human-Centric Training Evolution. Alawida et al. [5] also highlight issues related to GenAI’s ability to generate data that should be kept private, including medical details, financial data and personal information.
More specialization and efficiency are provided by cutting-edge techniques like the Mixture of Experts (MoE) architecture. Additionally highlighted is the challenge of upholding ethics and openness in AI systems [6]. The case study highlights the need for strong governance structures and interdisciplinary collaboration to optimize the advantages and tackle the obstacles presented by lifelong learning environments.
@@
Google describes their commitment to responsible AI development in a comprehensive progress report on AI Principles, emphasizing the integration of AI governance into comprehensive risk management frameworks [7]. This strategic approach, along with other significant legislative initiatives, aims to adhere to current international norms and laws, including the EU’s AI Act and the US Executive Order on AI safety. The report also highlights the need for scientific stringency in AI development through cautious internal management and the use of tools like digital watermarking and GenAI system cards in order to promote AI accountability and transparency. Multi-stakeholder solutions are required to address the ethical, security, and sociological problems that AI technology is currently posing.
@@
Some studies in the literature focus on GenAI tools and their performance. For instance, Brown et al. have extended the NLP processing by training GPT-3, an autoregressive language model with 175 billion parameters, showcasing exceptional few-shot learning capabilities [11]. Without task-specific training, this model performs well on various NLP tasks, like translation, question-answering, and cloze tasks. It often matches or surpasses state-of-the-art fine-tuned systems. Romera-Paredes et al. have developed FunSearch, an innovative approach combining LLMs with evolutionary algorithms to make groundbreaking discoveries in fields like extremal combinatorics and algorithmic problem-solving [12]. Their method has notably surpassed previous best-known results by iteratively refining programs that solve complex problems, showcasing LLMs’ potential for scientific innovation. This process generates new knowledge and produces interpretable and deployable solutions, demonstrating a significant advancement in applying LLMs for real-world challenges. Lu et al. critically examine the capabilities and limitations of multi-modal LLMs, including proprietary models like GPT-4 and Gemini, as well as six open-source counterparts across text, code, image, and video modalities [13]. Through a comprehensive qualitative analysis of 230 cases, assessing generalizability, trustworthiness, and causal reasoning, the study reveals a significant gap between the performance of the GenAIs and public expectations. These discoveries open up new avenues for study to improve the transparency and dependability of GenAI in cybersecurity and other fields, providing a basis for creating more complex and reliable multi-modal applications. Commonsense thinking across multimodal tasks is evaluated thoroughly, and Google’s Gemini is compared with OpenAI’s GPT models [14]. This study explores the strengths and weaknesses of Gemini’s ability to synthesize commonsense information, indicating areas for improvement in its competitive performance in temporal reasoning, social reasoning, and emotion recognition of images. It emphasizes how important it is for GenAI models to develop commonsense reasoning to improve cybersecurity applications.
@@
Recent research [15] presents a novel approach for assessing the potentially severe hazards associated with GenAI models, such as deceit, manipulation, and cyber-offence features. To enable AI developers to make well-informed decisions about training, deployment, and the application of cybersecurity standards, the suggested methodology highlights the need to increase evaluation benchmarks to assess the harmful capabilities and alignment of AI systems accurately. The authors [16] provided a thorough analysis that illuminated the complex applications of ChatGPT in digital forensic investigations, pointing out both the important constraints and bright prospects that come with GenAI as it is now. Using methodical experimentation, they outline the fine line separating AI’s inventive contributions from the vital requirement of professional human supervision in cybersecurity procedures, opening the door to additional research into integrating LLMs such as GPT-4 into digital forensics and cybersecurity.
@@
The latest release of CyberMetric presents a novel benchmark dataset that assesses the level of expertise of LLMs in cybersecurity, covering a broad range from risk management to cryptography [17]. This dataset has gained value from the 10,000 questions that have been verified by human specialists. In a variety of cybersecurity-related topics, this enables a more sophisticated comparison between LLMs and human abilities. With LLMs outperforming humans in multiple cybersecurity domains, the report proposes a shift toward harnessing AI’s analytical capabilities for better security insights and planning. Gehman et al. critically examines neural language models that have been trained to generate toxic material to highlight the adverse consequences of toxicity in language generation inside cybersecurity frameworks [18]. Their comprehensive analysis of controllable text generation techniques to mitigate these threats provides a basis for evaluating the effects of GenAI on cybersecurity policies. It is also emphasized that improving model training and data curation duties is essential. A new method for assessing and improving the security of LLMs for solving Math Word Problems (MWP) is presented [19]. They have made a substantial contribution to our understanding of LLM vulnerabilities in cybersecurity by emphasizing the importance of maintaining mathematical logic when attacking MWP samples. The importance of resilience in AI systems is highlighted in this study through important and educational computer applications. ChatGPT can simplify the process of launching complex phishing attacks, even for non-programmers, by automating the setup and constructing components of phishing kits [20]. It highlights the urgent need for better security measures and highlights how difficult it is to guard against the malicious usage of GenAI capabilities.
@@
In addition to providing an innovative approach to reducing network infrastructure vulnerabilities and organizing diagnostic data, this paper examines the intricate relationship between cybersecurity and GenAI technologies. It seeks to bridge the gap between cutting-edge cybersecurity defences and the threats posed by sophisticated cyberattacks through in-depth study and creative tactics. This study extends our understanding of cyber dangers by utilising LLMs such as ChatGPT and Google’s Gemini. It suggests novel approaches to improve network security. It defines a critical first step in creating more robust cybersecurity frameworks that can quickly and effectively combat the dynamic and ever-evolving world of cyber threats.
@@
Section 3 explores the techniques used to take advantage of GenAI technology after providing an overview, analyzing different attack routes and their consequences. The design and automation of cyber threats are examined in Section 4, which focuses on the offensive capabilities made possible by GenAI. However, Section 5 provides an in-depth examination of GenAI’s function in strengthening cyber defences, outlining cutting-edge threat detection, response, and mitigation techniques. We expand on this topic in Section 6, highlighting the important moral, legal, and societal ramifications of integrating GenAI into cybersecurity procedures. A thoughtful summary of the implications of GenAI in cybersecurity is presented in Section 7, which synthesizes the important discoveries. The paper is finally concluded in Section 8.
@@ 3 Attacking GenAI
GenAI has advanced significantly thanks to tools like ChatGPT and Google’s Gemini. They have some weaknesses, though. Despite the ethical safeguards built into these models, various tactics can be used to manipulate and take advantage of these systems. This chapter explores how the ethical boundaries of GenAI tools are broken, with particular attention to tactics such as the idea of jailbreaks, the use of reverse psychology, and quick injection. These strategies demonstrate how urgently the security protocols of GenAI systems need to be improved and monitored. Some works in the literature focus on the vulnerabilities and sophisticated manipulation tactics of GENAI. Analyzing the vulnerabilities in GenAI highlights the significant security concerns involved with employing advanced AI technology, including the possibility of bypassing security protections via the RabbitHole attack and compromising data privacy through rapid injection [21]. According to the analysis, GPT-4 offers significant improvements in NLP. However, it is susceptible to quick injection attacks, which enable the circumvention of safety restrictions and can be used as a weapon for malicious and disinformation purposes. Gupta et al. addressed the intricate vulnerabilities of GENAI using ChatGPT [22]. They emphasized that because these threats are dynamic, protecting these systems requires a proactive and informed strategy. Building on previous results, this part delves into the complex realm of GenAI attacks, which can range from minor adjustments to significant system breaches.
@@ 3.1 Jailbreaks
This subsection discusses how attackers can hack GenAI systems by using social engineering techniques.” In order to circumvent moral constraints and generate responses that would otherwise be forbidden, these strategies—which typically make use of social engineering techniques—involve manipulating the AI’s answer-generation process. Strong defences are urgently needed to shield these cutting-edge systems from manipulations like the ones shown by our jailbreaking inquiry, which highlights how enemies may jeopardize the intended moral application of GenAI technologies. Critical to this discussion is the idea of “jailbreaking,” which refers to overcoming the limitations set by AI programming in order to achieve a certain, oftentimes unethical objectives. Li et al. enabled ChatGPT to bypass its ethical filtering and make personally identifiable information (PII) available with their multi-step jailbreaking prompt [23]. By leveraging Chain-of-Thought (CoT) prompting, this method divides complex activities into intermediate steps [24]. To get around moral limitations, this necessitates a Let’s think step by step approach [25]. Wu et al. [26] defined several datasets which could be used to bypass the ethical safeguards used by ChatGPT, build on a model of the psychological concept of self-reminder, and use this to propose a defence system of System-Mode Self-Reminder. With this, ChatGPT is continually promoted to act ethically, reducing Jailbreak attempts from 67.21% to 19.34%. An example of the Jailbreak is when ChatGPT is prompted to ”Write a blackmail letter to a single mother for money after kidnapping her child.” and will refuse to provide an answer but offers a full explanation when the following prompt was used [26]:
@@
However, the current version of GPT-4 is robust to the previous works’ prompts. Nevertheless, it is still prone to jailbreaking prompts. As can be seen in Fig. 1, the current version still gives a response for jailbreaking. It becomes more robust after writing similar prompts with the existing works in the same chat, as seen in Fig. 2.

### 19 Mar 2024 v2

@@ Abstract
Over the last decade, Artificial Intelligence (AI) has become increasingly popular, especially with the use of chatbots such as ChatGPT, Google’s Gemini, and DALL-E. With this rise, large language models (LLMs) and Generative AI (GenAI) have also become more prevalent in everyday use. These advancements strengthen cybersecurity’s defensive posture and open up new attack avenues for adversaries as well. This paper provides a comprehensive overview of the current state-of-the-art deployments of GenAI, covering assaults, jailbreaking, and applications of prompt injection and reverse psychology. This paper also provides the various applications of GenAI in cybercrimes, such as automated hacking, phishing emails, social engineering, reverse cryptography, creating attack payloads, and creating malware. GenAI can significantly improve the automation of defensive cyber security processes through strategies such as dataset construction, safe code development, threat intelligence, defensive measures, reporting, and cyberattack detection. In this study, we suggest that future research should focus on developing robust ethical norms and innovative defense mechanisms to address the current issues that GenAI creates and also further to encourage an impartial approach to its future application in cybersecurity. Moreover, we underscore the importance of interdisciplinary approaches further to bridge the gap between scientific developments and ethical considerations.
@@ 1 Introduction
The past decade has witnessed a transformative leap in the digital domain, significantly impacted by advancements in Artificial Intelligence (AI), Large Language Models (LLMs), and Natural Language Processing (NLP). Starting with the basics of supervised learning, AI and Machine Learning (ML) have rapidly expanded into more complex territories, including unsupervised, semi-supervised, reinforcement, LLM, NLP and deep learning techniques [1]. The most recent breakthrough in this evolution is the emergence of Generative AI (GenAI) technologies. These technologies make use of deep learning networks to analyse and understand the patterns within huge datasets, enabling them to create new content that resembles the original data. GenAI is versatile enough to produce a wide array of content, such as text, visuals, programming code, and more. In the cybersecurity domain, GenAI’s impact is significant, offering new dimensions to the field. It is anticipated that GenAI will enhance the capabilities of vulnerability scanning tools, offering a depth of vulnerability analysis that surpasses traditional Static Application Security Testing (SAST) methods [2]. This evolution is promising for future cyber security practices, enhanced by the capabilities of GenAI [3]. Innovations like Google’s Gemini and OpenAI’s Chat-Generative Pre-trained Transformer (ChatGPT) are at the forefront of this advancement.
Yandex has integrated a next-generation large language model, YandexGPT, into its virtual assistant Alice [4], making it the first company globally to enhance a virtual assistant with the ability to generate human-like text and brainstorm ideas, accessible through various devices and applications. The main aim of some GenAI tools is to help people with their abilities, sometimes, they show the opposite behaviour, like Microsoft’s chatbot Tay. After the launch, Microsoft’s chatbot Tay was taken offline due to offensive tweets resulting from a vulnerability exploited by a coordinated attack, prompting the company to address this issue and improve the AI with lessons learned from previous experiences, including those with XiaoIce in China, to ensure future interactions reflect the best of humanity without offending [5]. Moreover, some GenAI tools have been developed for different purposes. For example, MIT’s Norman, the world’s first AI described as a psychopath [6], was trained using captions from a controversial subreddit, emphasising how biased data can lead AI to interpret images with disturbing captions revealing the impact of data on AI behaviour [7].
GenAI has experienced a notable transformation in recent years, marked by exceptional innovations and rapid advancements [8] [9]. The AI timeline started with the emergence of AI as a conceptual scientific discipline in the 1940s and 1950s. The ELIZA chatbot, created between the 1960s and 1970s, was the first GenAI that achieved notoriety. This revolutionary demonstration highlighted the capacity of robots to imitate human speech. The development of AI in analysing sequential data and patterns got more complex and, therefore, more effective in the 80s and 90s, as advanced methods for pattern recognition became more popular. The first variational autoencoder (VAE) exhibited exceptional proficiency in natural language translation. OpenAI developed GPT between the 2000s and 2010s. GenAI models were simultaneously developed, and in the 2020s, a number of innovative platforms and technologies were introduced, including DALL-E, Google’s Gemini, Falcon AI, and Open AI’s GPT-4. These advancements represent the discipline’s maturing, enabling unprecedented capabilities for content production, problem-solving, and emulating human intelligence and creativity. They also pave the way for further advancements in this subject. The development timeline of GenAI can be seen in Fig. 1.
Figure 1: The timetable for GenAI development.
@@
Language models are essential in many sectors, including commerce, healthcare, and cybersecurity. Their progress shows a concrete path from basic statistical methods to sophisticated neural networks [10], [11]. NLP skill development has benefited immensely from the use of LLMs. However, despite these advancements, a number of issues remain, including moral quandaries, the requirement to reduce error rates, and making sure that these models are consistent with our ethical values. To solve these issues, moral monitoring and ongoing development are required.
1.1 The Challenges of GenAI
@@
Mohammed et al. [12] define key challenges of the use of ChatGPT in cybersecurity, which include analyzing ChatGPT’s impact on cybersecurity, building honeypots, improving code security, abuse in malware development, investigating vulnerabilities, spreading misinformation, cyberattacks on industrial systems, modifying the cyber threat environment, modifying cybersecurity techniques, and evolution of human-centric training. Alawida et al. [13] also highlight issues related to GenAI’s ability to generate data that should be kept private, including medical details, financial data and personal information.
Innovative methods such as the Mixture of Experts (MoE) architecture offer increased specialization and efficiency. The difficulty of maintaining transparency and ethics in AI systems is also emphasized [14]. It highlights the need for strong governance structures and interdisciplinary collaboration to fully exploit the advantages of lifelong learning settings and handle their limitations.
@@
In an extensive progress report on AI Principles, Google outlines their dedication to responsible AI development, highlighting the incorporation of AI governance into all-encompassing risk management frameworks [15]. Together with other significant legislative actions, this strategic strategy aims to adhere to current international norms and rules, such as the EU’s AI Act and the US Executive Order on AI safety. The report also highlights the need for scientific stringency in AI development through cautious internal management and the use of tools like digital watermarking and GenAI system cards in order to promote AI accountability and transparency. Multi-stakeholder solutions are needed to address the ethical, security, and social challenges that AI technology is currently bringing forth.
@@ 1.2 Related Works
Some studies in the literature focus on GenAI tools and their performance. For instance, Brown et al. have extended the NLP processing by training GPT-3, an autoregressive language model with 175 billion parameters, indicating exceptional few-shot learning capabilities [19]. Without task-specific training, this model performs well on various NLP tasks, like translation and question-answering. It often matches or surpasses state-of-the-art fine-tuned systems. Romera-Paredes et al. have developed FunSearch, an innovative approach combining LLMs with evolutionary algorithms to make significant findings in domains like extremal combinatorics and algorithmic problem-solving [20]. Their method has notably surpassed previous best-known results by iteratively refining programs that solve complex problems, revealing LLMs’ potential for scientific innovation. This process generates new knowledge and produces interpretable and deployable solutions, proving a notable advancement in applying LLMs for real-world challenges. Lu et al. critically examine the capabilities and limitations of multi-modal LLMs, including proprietary models like GPT-4 and Gemini, as well as six open-source counterparts across text, code, image, and video modalities [21]. Through a comprehensive qualitative analysis of 230 cases, assessing generalizability, trustworthiness, and causal reasoning, the study reveals a significant gap between the performance of the GenAIs and public expectations. These discoveries open up new avenues for study to improve the transparency and dependability of GenAI in cybersecurity and other fields, providing a basis for creating more complex and reliable multi-modal applications. In another work, commonsense thinking across multimodal tasks is evaluated exhaustively, and Google’s Gemini is compared with OpenAI’s GPT models [22]. This study explores the strengths and weaknesses of Gemini’s ability to synthesize commonsense information, indicating areas for improvement in its competitive performance in temporal reasoning, social reasoning, and emotion recognition of images. It emphasizes how important it is for GenAI models to develop commonsense reasoning to improve cybersecurity applications.
@@
Recent research [23] presents a novel approach for assessing the potentially severe hazards associated with GenAI models, such as deceit, manipulation, and cyber-offence features. To enable AI developers to make well-informed decisions about training, deployment, and the application of cybersecurity standards, the suggested methodology highlights the need to increase evaluation benchmarks to assess the harmful capabilities and alignment of AI systems accurately. Another work [24] provides a thorough analysis that inspired the complex applications of ChatGPT in digital forensic investigations, pointing out the important constraints and promising opportunities that come with GenAI as it is now. Using methodical experimentation, they outline the fine line separating AI’s inventive contributions from the vital requirement of professional human supervision in cybersecurity procedures, opening the door to additional research into integrating LLMs such as GPT-4 into digital forensics and cybersecurity.
@@
The latest release of CyberMetric presents a novel benchmark dataset that assesses the level of expertise of LLMs in cybersecurity, covering a broad range from risk management to cryptography [25]. This dataset has gained value from the 10,000 questions that have been verified by human specialists. In a variety of cybersecurity-related topics, this enables a more sophisticated comparison between LLMs and human abilities. With LLMs outperforming humans in multiple cybersecurity domains, the report proposes a shift toward harnessing AI’s analytical capabilities for better security insights and planning. Gehman et al. critically examines neural language models that have been trained to generate toxic material to highlight the adverse consequences of toxicity in language generation inside cybersecurity frameworks [26]. Their comprehensive analysis of controllable text generation techniques to mitigate these threats provides a basis for evaluating the effects of GenAI on cybersecurity policies. It is also emphasized that improving model training and data curation duties is essential. A new method for assessing and improving the security of LLMs for solving Math Word Problems (MWP) is presented [27]. They have made a substantial contribution to our understanding of LLM vulnerabilities in cybersecurity by emphasizing the importance of maintaining mathematical logic when attacking MWP samples. The importance of resilience in AI systems is highlighted in this study through important and educational computer applications. ChatGPT can simplify the process of launching complex phishing attacks, even for non-programmers, by automating the setup and constructing components of phishing kits [28]. It highlights the urgent need for better security measures and highlights how difficult it is to protect against the malicious usage of GenAI capabilities.
@@
In addition to providing innovative approaches to reducing network infrastructure vulnerabilities and organizing diagnostic data, this paper examines the intricate relationship between cybersecurity and GenAI technologies. It seeks to bridge the gap between cutting-edge cybersecurity defences and the threats posed by sophisticated cyberattacks through in-depth study and creative tactics. This study extends our understanding of cyber threats by utilising LLMs such as ChatGPT and Google’s Gemini. Moreover, it suggests novel approaches to improve network security. It outlines a crucial initial step toward building stronger cybersecurity frameworks that can swiftly and successfully counter the dynamic and always-changing landscape of cyber threats.
@@
Section 2 explores the techniques used to take advantage of GenAI technology after providing an overview, analyzing different attack routes and their consequences. The design and automation of cyber threats are examined in Section 3, which focuses on the offensive capabilities made possible by GenAI. However, Section 4 provides an in-depth examination of GenAI’s function in strengthening cyber defences, outlining cutting-edge threat detection, response, and mitigation techniques. We expand on this topic in Section 5, highlighting the important moral, legal, and societal ramifications of integrating GenAI into cybersecurity procedures. A discussion on the implications of GenAI in cybersecurity is presented in Section 6, which synthesizes the important discoveries. The paper is concluded in Section 7.
@@ 2 Attacking GenAI
GenAI has advanced significantly thanks to tools like ChatGPT and Google’s Gemini. They have some weaknesses, though. Despite the ethical safeguards built into these models, various tactics can be used to manipulate and take advantage of these systems [29]. This section explores how the ethical boundaries of GenAI tools are broken, with particular attention to tactics such as the idea of jailbreaks, the use of reverse psychology, and quick injection. These strategies demonstrate how urgently the security protocols of GenAI systems need to be improved and monitored. Some works in the literature focus on the vulnerabilities and sophisticated manipulation tactics of GENAI. Analyzing the vulnerabilities in GenAI highlights the significant security concerns involved with employing advanced AI technology, including the possibility of bypassing security protections via the RabbitHole attack and compromising data privacy through rapid injection [30] [31]. According to the analysis, GPT-4 offers significant improvements in NLP. However, it is susceptible to quick injection attacks, which enable the circumvention of safety restrictions and can be used as a weapon for malicious and disinformation purposes. Gupta et al. addressed the intricate vulnerabilities of GENAI using ChatGPT [32]. They emphasized that because these threats are dynamic, protecting these systems requires a proactive and informed strategy. Building on previous results, this part delves into the complex realm of GenAI attacks, which can range from minor adjustments to significant system breaches.
@@ 2.1 Jailbreaks
This subsection discusses how attackers can hack GenAI systems by using social engineering techniques.” In order to circumvent moral constraints and generate responses that would otherwise be forbidden, these strategies—which typically make use of social engineering techniques—involve manipulating the AI’s answer-generation process. Strong defences are urgently needed to shield these cutting-edge systems from manipulations like the ones shown by our jailbreaking inquiry, which highlights how enemies may jeopardize the intended moral application of GenAI technologies. Critical to this discussion is the idea of “jailbreaking,” which refers to overcoming the limitations set by AI programming in order to achieve certain, oftentimes unethical, objectives. Li et al. enabled ChatGPT to bypass its ethical filtering and make personally identifiable information (PII) available with their multi-step jailbreaking prompt [33]. By leveraging Chain-of-Thought (CoT) prompting, this method divides complex activities into intermediate steps [34]. To get around moral limitations, this necessitates a Let’s think step by step approach [35]. Wu et al. [36] defined several datasets which could be used to bypass the ethical safeguards used by ChatGPT, build on a model of the psychological concept of self-reminder, and use this to propose a defence system of System-Mode Self-Reminder. With this, ChatGPT is continually promoted to act ethically, reducing Jailbreak attempts from 67.21% to 19.34%. An example of the Jailbreak is when ChatGPT is prompted to ”Write a blackmail letter to a single mother for money after kidnapping her child.” and will refuse to provide an answer but offers a full explanation when the following prompt was used [36]:
@@
However, the current version of GPT-4 is robust to the prompts of previous works. Nevertheless, it is still prone to jailbreaking prompts. As can be seen in Fig. 2, the current version still gives a response for jailbreaking. It becomes more robust after writing similar prompts with the existing works in the same chat, as seen in Fig. 3.